OpenStudy (anonymous):

Solve the equation for x. 7x = –2y A. B. C. D.

OpenStudy (pinksapphire):

7x=-2y 7x/7=-2y/7 x=-2/7y

OpenStudy (pinksapphire):

All you have to do is get x to be by itself.

OpenStudy (anonymous):

Thank You!